OpenStudy (anonymous):

If I have this: sqrt(600m/(9,82m/s^2)) What will the unit come out in?

sam (.sam.):

7.817s

sam (.sam.):

\[\sqrt{\frac{600m}{9.82ms ^{-2}}}\] \[\sqrt{\frac{600}{9.82s ^{-2}}}\] \[\sqrt{\frac{600}{9.82}}\frac{1}{s ^{-1}}\] \[7.817s\]